OpenStudy (blaque678):

Find an equation of the line that satisfies the given conditions. y-intercept 2; parallel to the line 5x + 6y + 7 = 0

OpenStudy (anonymous):

@blues when two lines are parallel , their slopes are equal u have been given the equation of the line , first determine the slope from this equation using y = mx +c form

OpenStudy (anonymous):

equation of the line is 5x+6y+7 = 0 6y = -5x-7 \[y = -\frac{ 5 }{ 6 }x-\frac{ 7 }{ 6 }\] now this is of the form of y = mx+c where m is slope and c is y intercept now can u tell what is the slope of the line ?

OpenStudy (anonymous):

slope is -5/6

OpenStudy (blaque678):

I get confused when I have to put in back to =0

OpenStudy (blaque678):

so it would be -7/6=-5/6+7=0?

OpenStudy (mrnood):

in the equation of form y=mx+c m is the slope. You have calculated the slope of the parallel line - so that is the slope of your new line. c is the intercept on the y axis, which is given in the question So now you know the slope and intercept - so you can write the equation y = mx + c using the values you have established....

OpenStudy (blaque678):

still doesn't make any since sorry

OpenStudy (mrnood):

do you understand the general form of a straight line equation? y=mx+c 

OpenStudy (mrnood):

The lines above are parallel - so they have the same slope They have different intercepts - line 2 is simply line 1 "shifted upwards"

OpenStudy (mrnood):

Do you want the answer to your question?

OpenStudy (mrnood):

apparently not